### v2

# RAGBench: Explainable Benchmark for Retrieval-Augmented Generation Systems

Abstract
Retrieval-Augmented Generation (RAG) has become a standard architectural pattern for incorporating domain-specific knowledge into user-facing chat applications powered by Large Language Models (LLMs). RAG systems are characterized by (1) a document retriever that queries a domain-specific corpus for context information relevant to an input query, and (2) an LLM that generates a response based on the provided query and context. However, comprehensive evaluation of RAG systems remains a challenge due to the lack of unified evaluation criteria and annotated datasets. In response, we introduce RAGBench: the first comprehensive, large-scale RAG benchmark dataset of 100k examples. It covers five unique industry-specific domains and various RAG task types. RAGBench examples are sourced from industry corpora such as user manuals, making it particularly relevant for industry applications. Further, we formalize the TRACe evaluation framework: a set of explainable and actionable RAG evaluation metrics applicable across all RAG domains. We release the labeled dataset at https://huggingface.co/datasets/rungalileo/ragbench. RAGBench explainable labels facilitate holistic evaluation of RAG systems, enabling actionable feedback for continuous improvement of production applications. Thorough extensive benchmarking, we find that LLM-based RAG evaluation methods struggle to compete with a finetuned RoBERTa model on the RAG evaluation task. We identify areas where existing approaches fall short and propose the adoption of RAGBench with TRACe towards advancing the state of RAG evaluation systems.
**footnotetext: Equal Contributions
1 Introduction
Despite remarkable reasoning and conversational abilities, out-of-the-box pre-trained Large Language Models (LLMs) struggle to reason about out-of-domain, knowledge-intensive queries [20, 13]. In response, Retriever-Augmented Generation (RAG) systems [20, 19] are becoming increasingly popular in user-facing dialogue applications [34]. Generally, RAG systems comprise a retriever component that queries relevant documents from an in-domain corpus and a downstream LLM generator model that incorporates the retrieved documents along with the original user query to output an informed response. The additional context helps ground the LLM in factual information and has been shown to boost performance on knowledge-intensive tasks [20].
Still, when used in production settings, RAG systems are prone to hallucinations as the generator model struggles to retrieve relevant information from the context [1, 30, 7]. In the absence of a one-fits-all approach, application-specific RAG systems must be fine-tuned for optimal performance on domain-specific tasks. However, the choice of retriever and generator models for each application is complex and has serious implications on overall system quality and costs. With numerous commercial and open-source generative LLMs readily available [^1] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard and many variable parameters in the RAG system design (Figure 1), tuning an optimal system for a particular RAG application involves iterative evaluation of multiple configurations. This motivates the need for automated RAG evaluation solutions.
In response, automated RAG evaluation systems like RAGAS [9] and TruLens [36] have emerged. These systems adopt a zero-shot LLM prompt-based approach to predict a set of curated RAG evaluation metrics. However, the lack of unified RAG benchmarks makes it difficult to compare approaches against each other. Each new study designs a new dataset, often employing LLMs as generators and labelers [9, 32, 4], which renders them irreproducible. A few benchmarks like RGB [4], AttributionBench [22] and RAGTruth [40] have been proposed recently, but they are small in size and target a disjoint set of labels. The exact RAG evaluation criteria also vary from study to study. ARES [32] and RAGAS [9] define a context relevance metric to evaluate the quality of the retrieved documents, along with answer relevance and faithfulness to evaluate the quality of the generative model. However, others have explored other metrics like correctness [1] noise rejection and robustness [4], to name a few. Finally, most studies evaluate on small in-domain evaluation datasets that are specific to each new application [32, 33, 9, 1, 4], leaving cross-domain generalization an open question.
In this work we propose RAGBench: a comprehensive dataset for training and benchmarking RAG evaluation models. RAGBench comprises data sourced from multiple domains along with a comprehensive suite of evaluation metrics. Specifically, we adopt existing metric definitions for context relevance, answer faithfulness [9, 32] and introduce two new metrics: context utilization and answer completeness. We argue that this new suite of metrics better describes the overall RAG system performance, with the potential to provide granular, actionable insights to the RAG practitioner.
We evaluate state-of-the art LLMs and existing RAG evaluation systems on RAGBench. We find that a 400M-parameter DeBERTa-large model that was fine-tuned on RAGBench outperforms few-shot LLM judges across numerous domains and task types. We highlight this result to motivate future work aimed at leveraging these data for advancing RAG evaluation models and improving on the proposed benchmark.
2 Related Work
We differentiate our work from existing ground-truth RAG datasets like ChatRAGBench [23], CRAG [41], DomainRAG [38], and ALCE [10]. These datasets contain RAG samples with ground-truth responses and are used for end-to-end evaluation of RAG systems via response-level metrics like exact match or ROUGE scores. In contrast, we design RAGBench to enable development of more mature evaluation systems that effectively evaluate different parts of the RAG system along multiple dimensions like retriever relevance, adherence and completeness of the response.
Various RAG benchmarks focus specifically on hallucination detection [40, 33, 21, 5]. FELM [5] is a multi-domain and task dataset with factuality labels for open domain QA. RAGTruth [40], DelucionQA [33], and HaluEval [21] are RAG-specific datasets with both synthetic as well as human-annotated labels for hallucinations in LLM reseponses. While these are appropriate benchmarks for hallucination detection models, they do not offer the level of granularity we offer with RAGBench that is necessary to understand the RAG system as a whole.
RAG evaluation
Recently, several parallel efforts have proposed approaches to automated RAG evaluation. In RAGAS [9], the authors query an LLM-judge (GPT-3.5) with a curated prompt to evaluate context relevance, answer relevance and faithfulness of a RAG response. Next, Saad-Falcon et al. [32] propose ARES, a framework for fine-tuning smaller NLI models to predict the same metrics. In parallel, Chen et al. [4] develop a heuristic system to probe LLM’s robustness to noisy and irrelevant context documents, and Adlakha et al. [1] explore heuristic algorithms to estimate RAG correctness and faithfulness. The lack of established RAG benchmarks makes it difficult to compare these approaches against each other. We aim to address this limitation by introducing RAGBench.
Finetuned RAG evaluation models
Fine-tuned LLM judges are another a common way to approach the LLM evaluation task [16, 44, 40]. A number of studies also leverage small, fine-tuned Natural Language Inference (NLI) models for RAG hallucination detection [2, 22, 32]. NLI models measure the degree of entailment between a premise and a hypothesis, which has been successfully repurposed for evaluating LLM response attribution in RAG setting. In this work, we train and evaluate an NLI model for RAG evaluation using RAGBench. The fine-tuned model not only outperforms LLM judges in hallucination/attribution detection but also excels on the new RAG evaluation metrics we propose.
Figure 1: RAG system workflow, with highlighted variable parameters: (1) Context format and length, (2) retriever model, (3) number of retrieved documents, and (4) generation model.
3 RAGBench Construction
3.1 Component Datasets
RAGBench is a collection of real-world datasets that span different domains and RAG task types. We source data from open-book Question-Answer (QA) datasets (CovidQA [26], PubmedQA [14], HotpotQA [42], MS Marco [28], CUAD [12], EManual [27], TechQA [3], FinQA [6], TAT-QA [47], ExpertQA [25], HAGRID [15]), as well one that was specifically adapted for RAG (DelucionQA [33]). We transform all 12 component datasets to a standardized RAG format with consistent annotations. To best represent real-world RAG scenarios, we vary a number parameters to construct the benchmark: the source domain, number of context documents, context token length, and the response generator model Figure 1 illustrates where these variable parameters fall in the RAG pipeline.
Source Domains
RAGBench comprises five distinct domains: bio-medical research (PubmedQA, CovidQA), general knowledge (HotpotQA, MS Marco, HAGRID, ExperQA), legal contracts (CuAD), customer support (DelucionQA, EManual, TechQA), and finance (FinBench, TAT-QA). We select these specific domains based on availability of data, and applicability to real-world RAG applications across different industry verticals. For detailed descriptions of each component data source, refer to Appendix 7.2.
Context Token Length
Context token length in RAGBench ranges from 100 to 11k tokens, which we report in Table 1. Notably, CUAD documents feature long contexts of up to 11k tokens each, compared to the relatively short context in PubMedQA.
Table 1: RAGBench component datasets.
Dataset Domain Document Source Question Source #docs doc length #Train #Dev #Test
PubMedQA biomedical
research
research
abstracts
automated heuristics 4 99 19.5k 2.5k 2.5k
CovidQA-RAG biomedical
research
research
papers
expert 4 122 2.5k 534 492
HotpotQA general
knowledge
wikipedia crowd-sourced 4 126 3.7k 847 776
MS Marco general
knowledge
web pages user
web queries
10 94 3.7k 790 839
HAGRID general knowledge wikipedia expert 3 153 2.0k 322 1.3k
ExpertQA general knowledge google search expert 3 548 1.6k 202 203
CUAD legal legal
contracts
expert 1 11k 1.5k 506 508
DelucionQA customer
support
Jeep manual LLM 3 296 1.5k 177 182
EManual customer
support
TV manual annotator 3 165 1k 132 132
TechQA customer
support
Technotes tech forums 5 1.8k 1.2k 302 310
FinQA finance earning
reports
expert 3 310 12k 1.7k 2.2k
TAT-QA finance financial
reports
expert 5 96 26k 3.2k 3.2k
Total 78k 12k 11k
Task Types
We curate RAGBench to inlcude a variety of difficult RAG task types. Customer support datasets simulate a common application of RAG in industry settings. FinQA and TAT-QA require numerical reasoning over hybrid tabular and text data. HotpotQA, CovidQA, and PubMedQA necessitate retrieval and reasoning over multiple context docs. The CUAD dataset is a challenging addition to RAGBench for several reasons: (i) it represents a difficult and highly-specialized real-world domain in which of-the-shelf pre-trained LLM models struggle to perform well [24], and (ii) it is equally challenging in RAG context due to very long context lengths of legal contract documents.
Question Sources
All component datasets include domain-specific questions that represent real-world user queries about various topics. Questions for DelucionQA, HotpotQA, and EManual are crowd-sourced; questions for CovidQA, CUAD, HAGRID, ExpertQA, and FinQA are composed by domain experts; MS Marco is sourced from real-world user web search queries; likewise, TechQA questions are user queries posted on IBM technical forums; PubMedQA is the only dataset with automatically-generated questions from research article titles.
Response Generation
For each component dataset we generate responses with LLMs. Exceptions to this are HAGRID and ExpertQA datasets, which contain LLM-generated responses in the original data. To introduce variability into the dataset, we generate two responses per input with different modes: GPT-3.5 (gpt-3.5-0125) and Claude 3 Haiku. Both are proprietary models that are offered at a reasonable price point [^2] https://openai.com/api/pricing/, https://www.anthropic.com/api, which we believe make them suitable candidates for generating real-world RAG responses. For CUAD we only generate responses with Claude 3 Haiku due to prohibitively long context lengths that exceed the GPT-3.5 16k token limit. To encourage a diverse distribution of labels in RAGBench, we use a basic prompt (Appendix 7.3) that does not explicitly require the model to stick to the provided context when generating the response. We set the temperature to 1.0 for generation.
Data Splits
We split each component dataset into train, validation, and test sets, ensuring there is no overlap in queries across splits from the same data source. RAGBench totals 100k samples, split across train, validation, and test sets. Component dataset statistics are reported in Table 1.
3.2 TRACe Evaluation Framework
We propose a suite of four comprehensive metrics to evaluate the quality of the retriever and the response generator components of RAG. An optimal RAG system must balance accuracy and efficiency. The retriever should precisely return all the necessary information to address the user query, avoiding any superfluous data. The generator must effectively utilize the retrieved information, ensuring the response is strictly based on the provided context without introducing any hallucinations in the output.
Towards comprehensive evaluation of the abovementioned criteria, we introduce the TRACe evaluation framework to measure uTilization, Relevance, Adherence, and Completeness of a RAG system. Utilization, Adherence, and Completeness measure the quality of the generator. Adherence here is synonymous with previously proposed answer faithfullness, groundednes, and attribution, all terms used in literature to measure how well an LLM output adheres to a source of factual information. Relevance measures the quality of the retriever output with respect to the query. Below we formalize the definition of each metric.
Definitions
Let $D$ be a set of context documents $\{d_{1}...d_{n}\}$ retrieved for a RAG input query. We define a set of relevant tokens in $d_{i}$ as $R_{i}=\{t_{1},...t_{r}\}$. $R_{i}$ encodes information in context document $d_{i}$ that is useful for answering the query. Similarly, we define $U_{i}=\{t_{1},...t_{u}\}$ as the set of utilized tokens in document $d_{i}$, which reflect information that the generation model is using to produce a response. Refer to Figure 2 for a visual representation of relevant and utilized spans. $Len(x)$ measures the length of strings in $x$, which can be interpreted as character length, token length, or sentence length. For calculating ground-truth metrics, we employ sentence-length, since it aligns best with our annotation schema (Section 3.4). However, token or character length may also be suitable for other use cases.
Figure 2: Example of RAG Question, Context, and Response. Relevant context spans are highlighted, and utilized spans are underlined.
Context Relevance
Context Relevance is defined in [9, 32] as the fraction of the retrieved context that is relevant to the input query. Low relevance points to an inefficient retriever that supplies excess information to the generation model. Long context inputs into the generator may accrue unnecessary costs, as well as compromise the quality of the generated output. We measure relevance of context document $d_{i}$ as:
document $\text{document relevance}=\frac{{Len(R_{i})}}{Len(d_{i})}$ (1)
Example-level relevance can be aggregated over all context documents in the example as:
example $\text{example relevance}=\frac{\sum_{i=1}^{|D|}{Len(R_{i})}}{\sum_{i=1}^{|D|}Len(d_{i})}$ (2)
Context Utilization
Context Utilization is a new metric introduced in TRACe. We aim to measure the the fraction of the retrieved context that is used by the generator to produce the response. Low Utilization in combination with low Relevance points to a greedy retriever, while low Utilization alone points to a weak generator that fails to leverage the provided context efficiently. Document-level and example-level Utilization are defined as:
$\text{document utilization}=\frac{{Len(U_{i})}}{Len(d_{i})}\quad\text{example utilization}=\frac{\sum_{i=1}^{|D|}{Len(U_{i})}}{\sum_{i=1}^{|D|}Len(d_{i})}$ (3)
Completeness
Completeness is another new metrics we introduce to measure how well the response incorporates all the relevant information in the context. Note that this is different from Utilization; it is possible to have high Relevance and high Utilization, but low Completeness when the generator utilizes irrelevant information in the context to produce a low quality response. Completeness for document $d_{i}$ is calculated as the fraction of utilized substrings among all relevant substrings:
$\text{completeness}=\frac{Len(R_{i}\cap U_{i})}{Len(R_{i})}$ (4)
And can be extended to example-level by considering all relevant and utilized substrings across all context documents.
Adherence
Adherence is designed to detect hallucinations in RAG responses. Our definition of Adherence is synonymous with answer faithfullness [9, 32], groundednes [36], and attribution [31]. For alignment with existing hallucination detection approaches, we define example-level adherence as a boolean indicating whether or not all parts of the response are grounded in the context. However, in our annotation schema (Section 3.4) we also define $A_{i}=\{t_{1},...t_{a}\}$ as the set of response tokens that are supported by the context to enable granular Adherence evaluation.
Figure 3: Distributions of relevance, utilization, and completeness labels in RAGBench. Y-axis is normalized to visualize densities.
3.3 RAGBench Statistics
RAGBench component datasets contain between 1% - 20% hallucinations. ExpertQA, CovidQA, and MS Marco contain the highest fraction of hallucinated responses (12%, 16%, and 13%, respectively), while Cuad, FinQA, and TAT-QA contain the least (about 1% for each). We visualize distributions of relevance, utilization, and completeness scores in Figure 3.
3.4 LLM annotator
We prompt GPT-4 (gpt-4-0125-preview) to produce ground truth Adherence, Relevance, and Utilization labels for input (documents, query, response) tuples in RAGBench. Completeness is easily derived from span-level Relevance and Utilization annotations, thus we don’t request explicit annotations for it.
For high quality labels, we use proven techniques like chain of thought [39] that have been shown to maximize the correlation between GPT-4 and human judgements [43, 46]. For relevance and utilization we request the LLM-annotator to directly identify relevant and utilized sub-strings in the input documents. For adherence, we instruct the LLM to identify which response sentences, if any, are supported by the provided context. We can then derive an example-level boolean adherence label by checking if all response sentences are supported. The exact prompt used for annotation is provided in Appendix 7.4. We apply post-processing steps to ensure high quality, reliable annotations from our GPT-labeler, which we outline in Appendix 7.5.
Table 2: GPT-4 annotator achieves high alignment with human judgements. We report F1 and Accuracy alignment metrics on human annotated subsets of DelucionQA. Adherence annotations are evaluated against DelucionQA human-annotated labels. Relevance and Utilization are evaluated against DelucionQA(40): a subset of 40 examples randomly sampled from the DelucionQA test set and annotated by the authors.
Test Set Metric F1 Accuracy
DelucionQA - example level Adherence 0.96 0.93
DelucionQA - span level Adherence 0.97 0.95
DelucionQA(40) - span level Utilization 0.92 0.94
DelucionQA(40) - span level Relevance 0.76 0.78
Alignment with Human Judgements
We validate our metric formulations and labeling approach on a human annotated benchmark. DelucionQA [33] is a curated collection of user queries on the operation of Jeep’s 2023 Gladiator model. Natural language queries are first generated by an LLM, then reviewed and filtered by human annotators. Context documents are sourced from Jeep’s Gladiator User Manual, and responses are generated by various LLMs. Combined with example-level and span-level annotations for hallucination, DelucionQA represents a realistic distribution of real-world user queries and RAG responses. We find that our GPT annotator achieves 93% and 95% example- and span-level agreement with human judgements on the DelucionQA test split (Table 2).
To validate relevance and utilization annotations, we also annotate a small subset of DelucionQA with granular relevance and utilization labels. We refer to this subset as DelucionQA(40) in Table 2. Similar to adherence, we observe high correlation between relevance and utilization judgements from GPT-4 and humans. Details of the annotation process and additional validations are found in Appendix 7.6.
(a) Retriever vs. Relevance
(b) LLM vs. Hallucination
(c) Prompt vs. Utilization
(d) Prompt vs. Completeness
Figure 4: Relationship between RAG system configuration and TRACe metrics. (a) The choice and configuration of RAG retriever component affects the average relevance of the retrieved context. In this example, a dense retriever with a low number of documents per query (k=2) yields the highest average context relevance. (b) The choice of LLM and generation prompt affect how well the RAG system utilizes the provided context. Prompting the LLM with a detailed chain-of-thought prompt leads to reduced hallucinations, as well as higher response utilization (c) and completeness (d) rates.
3.5 RAG Case Study
We design a case study to further motivate and validate the proposed TRACe framework. We sample 100 realistic world knowledge queries from the RAGTruth [40] QA training set, along with 2,512 unique context document chunks from the same dataset to use as inputs into our mock RAG systems. We simulate RAG setups of varying quality by controlling four configurable parameters: the retriever (sparse TF-IDF vs. dense), number of retrieved documents (2 vs. 4), generation model (GPT-3.5-turbo vs. GPT-4o), and 4 versions of the generation prompt template. For illustrative purposes, we evaluate one prompt template comprising of only the question (no context) vs. three RAG-style templates that encourage the LLM to utilize the provided context and/or utilize chain-of-thought (CoT). The full generation templates are provided in Appendix 7.7. We generate responses for the 100 RAGTruth queries with each of the 32 resulting RAG systems and use our LLM annotation prompt to evaluate the TRACe metrics.
Figure 4 demonstrates how the different RAG configurations influence TRACe metrics. For example, we confirm that the choice of the generative LLM model affects the amount of hallucinations (or non-adherent responses) in the generated text. Not surprisingly, GPT-4o combined with a chain-of-thought prompt yields the lowest amount of hallucination compared to the weaker GPT-3.5 model and less detailed prompts. Curiously, we find that GPT-3.5 hallucinates more when prompted with CoT, which may point to limitations in the model’s ability to reason about complex concepts 4(b). Overall, we find that prompting the LLM to think step-by-step and explain its reasoning leads to higher utilization of the provided context and more complete responses (4(c), 4(d)). Finally, we show that the choice of the retriever affects average relevance of the retrieved context documents per query 4(a).
4 Experiments
4.1 LLM Judge
We benchmarks a few LLM evaluators on RAGBench: (1) zero-shot GPT-3.5-judge, where we query GPT-3.5 with our annotation prompt, (2) RAGAS [9], and (3) TruLens [36]. RAGAS employs a series of few-shot prompts to GPT-3.5 to measure answer groundedness (Adherence) and Context Relevance metrics. Trulens is another zero-shot prompting approach that measures answer faithfulness (Adherence) and Context Relevance.
4.2 Fine-tuned Judge
We fine-tune a DeBERTa-v3-Large [11] NLI checkpoint [^3] https://huggingface.co/MoritzLaurer/DeBERTa-v3-large-mnli-fever-anli-ling-wanli from Laurer et al. [18] with one key architecture modification: we add a shallow prediction head for each of the output RAG metrics, which allows us to compute all TRACe metrics in a single forward pass. This is both cost-effective and enables transfer learning from head to head through back-propagation down to the shared base layers. Each prediction head is a single layer feed-forward net that acts on the token-level output of the last DeBERTa layer.
We attach two heads on the context tokens to estimate Relevance and Utilization probabilities, and another head on the response tokens to estimate Adherence. For training, we broadcast sentence-level annotations to tokens, and tune to maximize token-level probabilities of Relevant, Utilized, and Adherent spans. At inference, we impose a probability threshold=0.5 to predict Relevant and Utilized spans and Adherent spans and calculate TRACe metrics using equations 2, 3, and 4. For comparison with existing hallucination detection approaches, we also aggregate Adherence probabilities across the entire response to produce an example-level response adherence label. For details about training and hyperparameters, refer to Appendix 7.8.
4.3 Evaluation
Our granular annotation schema allows for various evaluation setups. For example, we could evaluate either span-level or example/response-level predictions. For easy comparison with existing RAG evaluation approaches that are less granular, we report area under the receiver-operator curve (AUROC) on the response-level hallucination detection task, and root mean squared error (RMSE) for example-level context Relevance and Utilization predictions.
Table 3: Benchmark evaluation on test splits. Reporting AUROC for predicting hallucinated responses (Hal), RMSE for predicting Context Relevance (Rel) and utilization (Util). indicates statistical significance at 95% confidence intervals, measured by bootstrap comparing the top and second-best results. RAGAS and Trulens do not evaluate Utilization.
GPT-3.5 RAGAS TruLens DeBERTA
[…]
PubMedQA 0.51 0.21 0.16 0.54 0.37 - 0.62 0.45 - 0.80 0.26 0.17
CovidQA-RAG 0.57 0.18 0.11 0.58 0.17 - 0.62 0.58 - 0.77 0.19 0.11
HotpotQA 0.59 0.11 0.08 0.62 0.14 - 0.64 0.73 - 0.85 0.11 0.08
MS Marco 0.65 0.23 0.11 0.63 0.25 - 0.62 0.61 - 0.70 0.22 0.10
HAGRID 0.58 0.22 0.15 0.62 0.22 - 0.67 0.69 - 0.81 0.20 0.13
ExpertQA 0.55 0.31 0.23 0.57 0.28 - 0.70 0.60 - 0.87 0.18 0.11
DelucionQA 0.57 0.18 0.10 0.70 0.22 - 0.55 0.64 - 0.64 0.15 0.10
EManual 0.54 0.17 0.11 0.57 0.27 - 0.61 0.64 - 0.76 0.13 0.13
TechQA 0.51 0.10 0.05 0.52 0.12 - 0.57 0.70 - 0.86 0.08 0.04
FinQA 0.57 0.10 0.13 0.57 0.06 - 0.53 0.79 - 0.81 0.10 0.10
TAT-QA 0.52 0.20 0.17 0.63 0.18 - 0.59 0.72 - 0.83 0.27 0.23
CUAD 0.51 0.27 0.11 0.66 0.19 - 0.40 0.66 - 0.80 0.24 0.10
5 Results
Table 3 reports results on test splits of each RAGBench component dataset. We compare baseline LLM methods with a finetunes DeBERTA encoder that trained on the full RAGBench train split.
We observe that the finetuned DeBERTa model trained on RAGBench achieves competitive performance with billion-parameter LLM judges across numerous domain subsets of the RAGBench test set. On the hallucination detection task, DeBERTA AUROC scores range from 0.64 to 0.86. While RMSE for relevance and utilization range from 0.04 to 0.26, depending on the domain and task.
Estimating Context Relevance is Difficult
As shown in Table 3, Relevance RMSE scores are generally higher than those for Utilization, indicating a greater difficulty in the relevance prediction task. Utilization can be assessed through a straightforward semantic comparison between the context and the response. In contrast, relevance is a more intricate metric. Due to the nature of RAG, the majority of retrieved documents are semantically related to the query. However, mere semantic similarity is insufficient. The model must ascertain whether the provided context includes specific information necessary to accurately answer the question. Thus, the task inherently involves deriving the correct answer, followed by assessing what information in the context may be used to arrive at that answer.
6 Conclusion
In this paper we introduce RAGBench, a large-scale dataset composed of real-world RAG examples intended for training and benchmarking powerful RAG evaluation models. To this end, we formulate TRACe, a RAG evaluation framework comprising four metrics: uTilization, Relevance, Adherence, and Completeness. TRACe standardizes the evaluation process, offering a consistent and systematic approach to measuring RAG system performance across various dimensions. We propose an automated approach to generate TRACe labels for RAGBench with an LLM and demonstrate high correlation between our approach and human judgements.
We benchmark existing RAG evaluation frameworks on RAGBench and demonstrate that a 400M parameter DeBERTa model finetuned on RAGBench data performs competitively with billion-parameter LLM Judges and commercial RAG evaluation systems. Though, the gap between the best-performing RAG evaluator and ground truth is still large. We motivate future work to leverage RAGBench toward fine-tuning more powerful evaluation models to explore the potential for narrowing the performance gap between these models and the ground truth.
Our contributions address the need for standardized benchmarks and methodologies, enabling more precise and actionable insights into the strengths and weaknesses of different RAG systems. This, in turn, will facilitate iterative improvement of RAG models, driving forward the capabilities of retrieval-augmented generation in real-world applications.